A square metal plate with a length of 0.20 m and with a charge of 9.7 x 10^-7 C. Find the electric field of the square metal plate.


1
Expert's answer
2021-03-31T09:31:47-0400
"E=\frac{\sigma}{\epsilon_0}=\frac{q}{2a^2\epsilon_0}"

"E=\frac{9.7\cdot10^{-7}}{2(0.2)^2(8.85\cdot10^{-12})}=1.37\cdot10^{6}\frac{N}{C}"